OpenStudy (anonymous):

Multiply and Simplify: (x + 4)(x + 3)

OpenStudy (anonymous):

2xy(7z + 12y)

OpenStudy (anonymous):

that is not right

OpenStudy (anonymous):

12x^2

OpenStudy (anonymous):

\[\huge Use \ (x+a) (x+b)=x^2+ax+bx+ab \\ \huge = x^2 + 4x + 3x + 12 \]

OpenStudy (anonymous):

@rannsan do u get that

OpenStudy (aripotta):

that's FOIL, right? i think it would be better if you explained what exactly to do rather than giving an equation, but that's just me

OpenStudy (anonymous):

I did not get that

OpenStudy (aripotta):

you have to use FOIL. First times First, Outer times Outer, Inner times Inner, and Last times Last

OpenStudy (anonymous):

1 multiply the first numbers in both the parenthesis x & x 2 multiply the outer numbers in both parenthesis ( x & b) 3 multiply the inner numbers (x&a) 4 multiply the last numbers ( a &b) Source(s): wiki

OpenStudy (anonymous):

ok this is gonna sound really stupid on my part but how can I multiply if I don't know what x is?

OpenStudy (aripotta):

 x times x equals 2x